Zvinakis Mining Company paid $290,000 for the rights to mine lead in southeast Missouri. The cost to drill and erect a mine shaft was $2,490,000, and equipment to process the lead ore before shipment to the smelter was $1,989,000. The mine is expected to yield 2,000,000 tons of ore during the five years it is expected to be operating. The equipment has an estimated residual value of $159,000 when mining is concluded. The mine started operations on April 30, 2018. In 2018, 390,000 tons of ore were extracted, and in 2019, 790,000 tons were mined.

Required:
1. Compute the depletion rate and the units-of-production depreciation rate.
2. Compute depletion and depreciation for 2018 and 2019.

Compute the depletion rate and the units-of-production depreciation rate.

Depletion rate per ton:

Depreciation rate per ton:

Compute depletion and depreciation for 2018 and 2019.

depletion 2018

depletion 2019

depreciation 2018

depreciation 2019
Answer #1
1
Mineral rights 290000
Add: Mine shaft 2490000
Depletion base 2780000
Divide by Recoverable ore 2000000
Depletion rate per ton 1.390
Mining equipment 1989000
Less: residual value 159000
Depreciable base 1830000
Divide by Recoverable ore 2000000
Depreciation rate per ton 0.915
2
Depletion 2018 542100 =390000*1.39
Depletion 2019 1098100 =790000*1.39
Depreciation 2018 356850 =390000*0.915
Depreciation 2019 722850 =790000*0.915